OpenStudy (anonymous):

What are the roots of the quadratic equation x2 − 2x + 4 = 8x – 21?

OpenStudy (help!!!!):

factor out or use quadratic equation

OpenStudy (bradely):

x^2 − 2x + 4 = 8x – 21? x^2 − 10x +25 =0 (x-5)^2=0 x=5,5 Source: http://www.mathskey.com/question2answer/